OpenStudy (anonymous):

FIND F(-1) and F(2) f(x)=sqrtx+14 f(-1)= f(2)=

OpenStudy (anonymous):

i take it \[F(x)=\int f(x)dx\] yes?

OpenStudy (anonymous):

is it \[f(x)=\sqrt{x+14}\] or \[f(x)=\sqrt{x}+14\]

OpenStudy (anonymous):

SECOND ONE

OpenStudy (anonymous):

SORRY....FIRST ONE!!!

OpenStudy (anonymous):

whew then it is easy \[F(x)=\frac{2}{3}x^{\frac{3}{2}}+14x\]

OpenStudy (anonymous):

still easy enough it is \[F(x)=\frac{2}{3}(x+14)^{\frac{3}{2}}\] \[F(x)=\frac{2}{3}\sqrt{(x+14)^3}\]